OpenStudy (anonymous):

What is the focus of the parabola y =1/8(x – 2)^2 + 4 ?

OpenStudy (anonymous):

the equation can be solved by considering z=1/y. so u have z=8(x-2)^2+4. Now the focus of this parabola(of the standard form ax^2 +bx +c) is \[(-b/2a, -b^{2}/4a + c +1/4a)\]. So now u have that focus....1/f should give u the coordinates of the original parabola.....

OpenStudy (anonymous):

huh? what is the focus?

OpenStudy (anonymous):

ok...see the euation becomes z=8x^2-32x+36 when u expand right ? so that gives a=8, b=-32, and c=36 in the general form. so the focus of the parabola z is (-2, 129/32). The inverse of this point is the focus of the parabola y. which is(-1/2,32/129).

OpenStudy (anonymous):

do u have the answer??

OpenStudy (anonymous):

no i dont i have no clue what u jst said these are my options: (2, –2) (–2, –2) (–2, 6) (2, 6)

OpenStudy (anonymous):

hmm, i think i might made a mistake in assuming the inverse operation. nevertheless, if u gave me those options, id go for (-2, 6). coz 2 cannot be the x coordinate. its surely -2. n 6 is more likely to be the y coordinate

OpenStudy (anonymous):

okay thank you :)